# An Upper Bound on the Error Probability of RPA Decoding of Reed-Muller Codes Over the BSC

### v1

@@ I Introduction
Reed-Muller (RM) codes are a family of binary linear codes that are obtained by the evaluations of Boolean polynomials on the points of the Boolean hypercube [1, 2]. They are of widespread interest in the information theory and error-control coding communities, owing to recent breakthrough theoretical progress has shown that RM codes by themselves are capacity-achieving for the binary erasure channel [3], and more generally, for BMS channels [4, 5]. Further, they share a close relationship [6, 7] with polar codes [8]—the first explicit family of (linear) codes that achieve capacity over binary-input memoryless symmetric (BMS) channels and also provably achieve high rates over many channels with memory [9, 10].
@@
In the light of their importance, much work has been dedicated to finding good decoding algorithms for RM codes. The earliest such algorithm by Reed [1] is capable of correcting bit-flip errors up to half the minimum distance of the RM code. For the special case of first-order RM codes, a Fast Hadamard Transform-based (or FHT-based) decoder was designed in [11, 12], which is an efficient implementation of a maximum likelihood (ML) decoding procedure. For second-order RM codes, a decoding algorithm that works well for moderate blocklength $N$ ($N\leq 1024$) RM codes was discussed in [13, 14]. Further, recursive decoding algorithms for RM codes (with and without lists) with good performance at moderate blocklengths were described and analyzed in [15, 16, 17]. We also refer the reader to several recent data-driven and heuristic approaches for decoding RM codes [18, 19, 20, 21].
@@
A much more recent decoding algorithm, variants of which were shown to achieve near-ML performance at moderate blocklengths over the binary symmetric channel (BSC), is the Recursive Projection-Aggregation (RPA) decoder of Ye and Abbe [22]. Via extensive simulation studies, the work [22] showed that the RPA algorithm outperforms the decoding algorithms in [13, 14] for second-order RM codes and the decoding algorithms in [15, 16, 17], for selected moderate blocklength RM codes. Later works [23, 24] presented procedures for reducing the complexity of the RPA algorithm. We mention however that recent work [25] suggests that running the RPA decoder by employing an aggregation step that uses only a subset of the total number of subspaces of a fixed dimension could lead to poor performance, due to the decoder’s potential failure in correcting so-called “coset error patterns".
@@
While it is now well-accepted that the RPA algorithm (as described in [22]) demonstrates good performance in practice for low-rate (in particular, fixed-order) RM codes, a formal, theoretical analysis of the RPA decoder has been largely lacking in the literature. In this paper, we make progress towards this goal, by obtaining explicit upper bounds on the error probabilities of the RPA decoder, when the channel used is the BSC. Importantly, we prove that for sufficiently large blocklengths, with overwhelmingly high probability, the RPA decoder recovers the correct input codeword in RM codes of order that scales up to logarithmically in the parameter $m$, where $N=2^{m}$. This result can hence be viewed as supplementing the results in [26, 27], where it was shown that RM codes of order growing linearly in the parameter $m$, but with vanishing rate, have error probabilities that decay to zero in the large blocklength limit. Our result, although weaker in terms of the orders of the RM codes of interest, establishes that an explicit, low-complexity decoding algorithm (namely, the RPA algorithm) can in fact be used to achieve vanishing error probabilities.
@@
Our analysis is based on explicit upper bounds on the probability of incorrect maximum likelihood (ML) decoding of first-order RM codes, which is used as a subroutine in the RPA decoder. With such upper bounds in place, we then characterize upper bounds on the probabilities of error in the “aggregation" steps carried out by the RPA decoder, conditioned on the event that all the “projections" are decoded correctly. Interestingly, our results show that even when one restricts attention to the event that each of the recursive calls to the RPA decoder converges in just one iteration, it is possible to obtain non-trivial performance estimates, which guarantee vanishing error probabilities for RM orders that increase with $m$.
@@
An interesting line of future work will be to characterize the largest order of an RM code, as a function of the parameter $m$, which can be decoded via RPA decoding with vanishing error probability, in the limit as the blocklength increases to infinity. Such analysis will require a finer understanding of the correlations among the random variables that are recursed on over several iterations of the RPA decoder, when potentially several runs of each recursive call to the RPA decoder are run. Importantly, via such more sophisticated analysis, we shall be able to determine if RM codes can achieve positive rates over the BSC under RPA decoding.
@@ II-B Reed-Muller Codes
It is known that the evaluation vectors of all the distinct monomials in the variables $x_{1},\ldots,x_{m}$ are linearly independent over $\mathbb{F}_{2}$. Hence, RM$(m,r)$ has dimension $\binom{m}{\leq r}:=\sum_{i=0}^{r}{m\choose i}$. We then have that RM$(m,r)$ is a $\left[2^{m},\binom{m}{\leq r}\right]$ linear code. It is also known that RM$(m,r)$ has minimum Hamming distance ${d}_{\text{min}}(\text{RM}(m,r))=2^{m-r}$.
@@ II-C System Model and the RPA Decoder
In this subsection, we set up the communication system of interest and briefly recall the Recursive Projection-Aggregation (RPA) algorithm introduced in [22], for this setting.
@@ II-C2 The RPA Decoder
For any one-dimensional subspace $\mathbb{B}_{i}$ of $\{0,1\}^{m}$, $i\in[N-1]$, that contains the non-zero vector $\mathbf{b}_{i}$, we let $\{0,1\}^{m}/\mathbb{B}_{i}$ denote the collection of cosets of $\mathbb{B}_{i}$ in $\{0,1\}^{m}$. For any coset $T\in\{0,1\}^{m}/\mathbb{B}_{i}$ that contains the vector $\mathbf{z}$ (we sometimes write such a coset as $[\mathbf{z}+\mathbb{B}_{i}]$) and for the given received vector $\mathbf{Y}$, we define $\mathbf{Y}_{/\mathbb{B}_{i}}(T):=\mathbf{Y}_{\mathbf{z}}\oplus\mathbf{Y}_{\mathbf{z}\oplus\mathbf{b}_{i}}$ as the “projection" of the vector $\mathbf{Y}$ onto the coset $T$. Finally, we define the projection of $\mathbf{Y}$ onto the cosets of $\mathbb{B}_{i}$ as
$\mathbf{Y}_{/\mathbb{B}_{i}}:=\left(\mathbf{Y}_{/\mathbb{B}_{i}}(T):\ T\in\{0,1\}^{m}/\mathbb{B}_{i}\right),$ (1)
for some fixed ordering among cosets $T$. One can define a projection $\mathbf{c}_{/\mathbb{B}_{i}}$ analogously for the input codeword $\mathbf{c}$ too; via standard properties of Reed-Muller codes, it can be argued that $\mathbf{c}_{/\mathbb{B}_{i}}\in\text{RM}(m-1,r-1)$, if $\mathbf{c}\in\text{RM}(m,r)$. We now briefly recall the RPA decoder for second-order RM codes. The main decoding algorithm is shown as Algorithm 1, which in turn uses the “aggregation" step in Algorithm 2 as a subroutine. We observe that the RPA decoding procedure consists of two key steps: (i) decoding of first-order ($r=1$) RM codes via the Fast Hadamard Transform (FHT) decoder (see [12] and [29, Sec. V.B.1]), which is simply an efficient implementation of a maximum likelihood (ML) decoder for first-order RM codes, and (ii) an aggregation step that aggregates multiple noisy estimates of a single codeword symbol.
We mention that our presentation of the RPA decoding algorithm and the aggregation subroutine are specialized to the setting where one projects of the received vector $\mathbf{Y}$ using one-dimensional subspaces, as against the more general formulation in [22], which allows for the use of subspaces of higher dimension. As we shall see, projections via one-dimensional subspaces by themselves give rise to correct decoding, with very high probability, for sufficiently large blocklengths $N$.
Algorithm 1 The RPA decoding algorithm for RM$(m,r)$
@@
4:         Compute the projections $\mathbf{Y}_{/\mathbb{B}_{i}}$, $i\in[N-1]$ as in (1).
5:         if $r\neq 2$ then
6:              For each $i\in[N-1]$, set $\widehat{\mathbf{Y}}_{/\mathbb{B}_{i}}\leftarrow$ RPA($\mathbf{Y}_{/\mathbb{B}_{i}},m-1,r-1,M_{\text{iter}}$).
@@
8:              For each $i\in[N-1]$, decode $\mathbf{Y}_{/\mathbb{B}_{i}}$ to $\widehat{\mathbf{Y}}_{/\mathbb{B}_{i}}$ using the FHT-based decoder.
9:         Compute $\overline{\mathbf{Y}}\leftarrow$ Aggregate$(\mathbf{Y},\widehat{\mathbf{Y}}_{/\mathbb{B}_{1}},\ldots,\widehat{\mathbf{Y}}_{/\mathbb{B}_{N-1}})$.
@@
1:procedure Aggregate($\mathbf{Y},\widehat{\mathbf{Y}}_{/\mathbb{B}_{1}},\ldots,\widehat{\mathbf{Y}}_{/\mathbb{B}_{N-1}}$)
2:     Initialize the length-$N$ vector $\textsc{Flip}\leftarrow\mathbf{0}$.
@@
4:         Compute $\phi(\mathbf{x})=\sum_{i=1}^{N-1}\mathds{1}\{Y_{\mathbb{B}_{i}}([\mathbf{x}+\mathbb{B}_{i}])\neq\widehat{Y}_{\mathbb{B}_{i}}([\mathbf{x}+\mathbb{B}_{i}])\}$.
5:         Set Flip$(\mathbf{x})=1$, if $\phi(\mathbf{x})>\frac{N-1}{2}$.
@@ III Main Results
An important fact about the error probabilities above is [22, Prop. 2], which states that for all $m\geq 1,1\leq r\leq m$, we have that the probability of error over the BSC$(p)$, upon transmission of a codeword $\mathbf{c}\in\text{RM}(m,r)$ is independent of $\mathbf{c}$. This, in turn, implies that we can focus our analysis on the case when $\mathbf{c}=\mathbf{0}\in\{0,1\}^{N}$, with $P_{\text{err}}(\text{RM}(m,r))=P_{\text{err}}(\mathbf{0})$. In all the analysis that follows, we shall make the simplifying assumption that the input codeword is fixed to be the all-zeros codeword $\mathbf{0}$.
Before we state our main result, we will require some more notation. Let $\overline{p}:=\frac{1}{2}\cdot(1-(1-2p)^{2^{r-2}})$ and for any $\alpha\in(0,\frac{1}{2})$, let $\eta(\alpha):=\frac{1}{2}\cdot(1-4\alpha(1-\alpha))$. Note that $\eta(\alpha)<\frac{1}{2}$, for all $\alpha\in(0,\frac{1}{2})$. Our main result is encapsulated in the following theorem:
@@
For any $0<\epsilon<\eta(\overline{p})$, we have that for $r\geq 2$,
$P_{\text{err}}(\text{RM}(m,r))\leq 32N^{r+1}\cdot e^{-2^{-(r+1)}\cdot{N\epsilon^{2}}}.$
@@
$\rho_{m}(r,\epsilon):=m(r+1)-\frac{\delta\log_{2}(e)}{8}\cdot 2^{m-r}\cdot(1-2p)^{2^{r}}+5.$
@@
Note, however, that RM codes with $r\leq\log(cm)$ as above have rate $R_{m}:=\frac{\log|\text{RM}(m,r)|}{N}=\frac{{m\choose\leq r}}{N}$ such that $R_{m}\xrightarrow{m\to\infty}0$. In other words, via the results in this paper, we only obtain that RPA decoding achieves vanishing error probabilities, in the limit as the blocklength goes to infinity, for selected RM codes that have vanishing rate. An important question hence, as mentioned earlier, is whether RM codes can achieve positive rate (or even the capacity of the BSC$(p)$) under RPA decoding—a fertile topic for future research.
Furthermore, for fixed (constant) orders $r$, Theorem III.1 also allows us to understand how quickly, as a function of $m$, the crossover probability $p$ (also called the “decoding radius") can converge to its maximum value that is $1/2$, while also ensuring that $P_{\text{err}}(\text{RM}(m,r))\xrightarrow{m\to\infty}0$. Observe that the statement that a decoding algorithm can achieve vanishing error probabilities, in the limit as the blocklength $N$ tends to infinity, is equivalent to the statement that the algorithm can correct $N(p-\delta)$ errors with high probability, for any $\delta\in(0,1)$. The following result hence establishes a lower bound on the number of such errors correctable with high probability, as a function of $m$, for fixed orders $r$, in a vein similar to the results in [15].
@@
This implies that the convergence of the decoding radius to $1/2$ (captured by $\gamma_{m}$), for fixed orders $r\geq 2$, under our analysis of RPA decoding, is slower than that obtainable using the decoding algorithm $\Phi_{r}^{m}$ in [15] (captured by $\beta_{m}$). However, for sufficiently large $m$, it can be verified similarly that the convergence of the decoding radius to $1/2$ under RPA decoding is faster than the decoding algorithm $\Psi_{r}^{m}$ in [15]. It is an interesting question to check whether the RPA decoder can outperform $\Phi_{r}^{m}$ in [15] in terms of the speed of convergence of the decoding radius, using more sophisticated analysis of the RPA error probabilities, possibly by considering cases where each RPA run converges after larger than one iteration, or by using higher dimensional subspaces for projection.
In Sections IV and V, we make progress towards proving III.1, formally. In the next section, we shall analyze the behaviour of the FHT decoder in Step 8 of Algorithm 1 above (which acts on the first-order RM codes obtained via the projection step) and the aggregation step in Step 9 of Algorithm 1, separately. Via such analysis, we shall obtain an understanding of the performance of the RPA decoder on second-order ($r=2$) RM codes. In Section V, we shall use the results obtained in the next section, recursively, to get estimates of the error probabilities of the RPA decoder for general orders $r>2$.

### v2

# An Analysis of RPA Decoding of Reed-Muller Codes Over the BSC

@@ I Introduction
Reed-Muller (RM) codes are a family of binary linear codes that are obtained by the evaluations of Boolean polynomials on the points of the Boolean hypercube [1, 2]. They are of widespread interest in the information theory and error-control coding communities, owing to recent breakthrough theoretical progress has shown that RM codes by themselves are capacity-achieving for the binary erasure channel [3], and more generally, for BMS channels [4, 5]. Further, they share a close relationship [6, 7] with polar codes [8]—the first explicit family of (linear) codes that was shown to achieve capacity over binary-input memoryless symmetric (BMS) channels and also provably achieve high rates over many channels with memory [9, 10].
@@
In the light of their importance, much work has been dedicated to finding good decoding algorithms for RM codes. The earliest such algorithm by Reed [1] is capable of correcting bit-flip errors up to half the minimum distance of the RM code. For the special case of first-order RM codes, a Fast Hadamard Transform-based (or FHT-based) decoder was designed in [11, 12], which is an efficient implementation of a maximum likelihood (ML) decoding procedure. For second-order RM codes, a decoding algorithm that works well for moderate blocklength $N$ ($N\leq 1024$) RM codes was discussed in [13, 14]. Further, recursive decoding algorithms for RM codes (with and without lists) that display good performance at moderate blocklengths were described and analyzed in [15, 16, 17, 18]. We also refer the reader to several recent data-driven and heuristic approaches for decoding RM codes [19, 20, 21, 22].
@@
A much more recent decoding algorithm, variants of which were shown to achieve near-ML performance at moderate blocklengths over the binary symmetric channel (BSC), is the Recursive Projection-Aggregation (RPA) decoder of Ye and Abbe [23]. Via extensive simulation studies, the work [23] showed that the RPA algorithm outperforms the decoding algorithms in [13, 14] for second-order RM codes and the decoding algorithms in [15, 16, 17], for selected moderate blocklength RM codes. Later works [24, 25] presented procedures for reducing the complexity of the RPA algorithm.
@@
While it is now well-accepted that the RPA algorithm (as described in [23]) demonstrates good performance in practice for low-rate (in particular, fixed-order) RM codes, a formal, theoretical analysis of the RPA decoder has been largely lacking in the literature. In this paper, we make progress towards this goal, by obtaining explicit upper bounds on the error probabilities of the RPA decoder, when the channel used is the BSC. Importantly, we prove that for sufficiently large blocklengths, with overwhelmingly high probability, the RPA decoder recovers the correct input codeword in RM codes of order that scales up to roughly logarithmically in the parameter $m$, where $N=2^{m}$. This result can hence be viewed as supplementing the results in [26, 27], where it was shown that RM codes of order growing linearly in the parameter $m$, but with vanishing rate, have error probabilities that decay to zero in the large blocklength limit. Our result, although weaker in terms of the orders of the RM codes of interest, establishes that an explicit, low-complexity decoding algorithm (namely, the RPA algorithm) can in fact be used to achieve vanishing error probabilities, for Reed-Muller orders that grow with $m$.
@@
Our analysis proceeds via first obtaining explicit upper bounds on the probability of incorrect maximum likelihood (ML) decoding of first-order RM codes, which is used as a subroutine in the RPA decoder. With such upper bounds in place, we then characterize upper bounds on the probabilities of error in the “aggregation" steps carried out by the RPA decoder, conditioned on the event that all the “projections" are decoded correctly. Interestingly, our results show that even when one restricts attention to the event that each of the recursive calls to the RPA decoder converges in just one iteration, it is possible to obtain non-trivial performance estimates.
@@
Our results can be seen as the first in a line of work to characterize the largest order of an RM code, as a function of the parameter $m$, which can be decoded via RPA decoding with vanishing error probability, in the limit as the blocklength increases to infinity. Such analysis, however, will require a finer understanding of the correlations among the random variables that are recursed on over several (more than one) iterations of the RPA decoder. Importantly, via such more sophisticated analysis, we shall be able to determine if RM codes can achieve positive rates over (and potentially the capacity of) the BSC under RPA decoding.
@@ II-B Reed-Muller Codes
It is known that the evaluation vectors of all the distinct monomials in the variables $x_{1},\ldots,x_{m}$ are linearly independent over $\mathbb{F}_{2}$. Hence, RM$(m,r)$ has dimension $\binom{m}{\leq r}:=\sum_{i=0}^{r}{m\choose i}$. We then have that RM$(m,r)$ is a $\left[2^{m},\binom{m}{\leq r}\right]$ linear code. It is also known that RM$(m,r)$ has minimum Hamming distance $d:={d}_{\text{min}}(\text{RM}(m,r))=2^{m-r}$.
@@ II-C System Model and the RPA Decoder
In this subsection, we set up the system model of interest and briefly recall the Recursive Projection-Aggregation (RPA) algorithm introduced in [23], for this setting.
@@ II-C2 The RPA Decoder
Fix a positive integer $k\leq r-1$, which denotes the dimension of the subspaces used in the RPA algorithm. We assume throughout this paper that $k$ divides $r-1$ (denoted as $k|(r-1)$). For any $k$-dimensional subspace $\mathbb{B}$ of $\{0,1\}^{m}$, $i\in[N-1]$, that contains the non-zero vector $\mathbf{b}$, we let $\{0,1\}^{m}/\mathbb{B}$ denote the collection of cosets of $\mathbb{B}$ in $\{0,1\}^{m}$. Now, consider the collection $\{\mathbb{B}_{i}\subseteq\{0,1\}^{m}\}$ of $k$-dimensional subspaces of $\{0,1\}^{m}$. The total number of such subspaces $n_{k,m}$ is given by the following Gaussian binomial coefficient [28, Ch. 15, Sec. 2, Thm. 9]:
[…]
Note that when $k=1$, we have $n_{1,m}=2^{m}-1=N-1$.
Given a set $\mathcal{S}\subseteq\{0,1\}^{m}$, we use the notation $\bigoplus_{\mathbf{x}\in\mathcal{S}}Y_{\mathbf{x}}$ to denote the modulo-$2$ sum of all coordinates of $\mathbf{Y}$ whose indices lie in $\mathcal{S}$. For any $i\in[n_{k,m}]$, any coset $T\in\{0,1\}^{m}/\mathbb{B}_{i}$ that contains the vector $\mathbf{x}$ (we denote such a coset as $[\mathbf{x}+\mathbb{B}_{i}]$), and for the given received vector $\mathbf{Y}$, we define ${Y}_{/\mathbb{B}_{i}}(T):=\bigoplus_{\mathbf{b}\in\mathbb{B}_{i}}Y_{\mathbf{x}\oplus\mathbf{b}}$ to be the “projection" of $\mathbf{Y}$ onto the coset $T\in\{0,1\}^{m}/\mathbb{B}_{i}$. Finally, we define the projection of $\mathbf{Y}$ onto the cosets of $\mathbb{B}_{i}$ as
$\mathbf{Y}_{/\mathbb{B}_{i}}:=\left({Y}_{/\mathbb{B}_{i}}(T):\ T\in\{0,1\}^{m}/\mathbb{B}_{i}\right),$ (2)
for some fixed ordering among cosets $T$. One can define a projection $\mathbf{c}_{/\mathbb{B}_{i}}$ analogously for the input codeword $\mathbf{c}$ too; via standard properties of Reed-Muller codes, it can be argued that $\mathbf{c}_{/\mathbb{B}_{i}}\in\text{RM}(m-k,r-k)$, if $\mathbf{c}\in\text{RM}(m,r)$. We now briefly recall the RPA decoder for second-order RM codes. The main decoding algorithm is shown as Algorithm 1, which in turn uses the “aggregation" step in Algorithm 2 as a subroutine. We observe that the RPA decoding procedure consists of two key steps: (i) decoding of first-order ($r=1$) RM codes via the Fast Hadamard Transform (FHT) decoder (see [12] and [29, Sec. V.B.1]), which is simply an efficient implementation of a maximum likelihood (ML) decoder for first-order RM codes, and (ii) an aggregation step that aggregates multiple noisy estimates of a single codeword symbol.
Algorithm 1 The RPA decoding algorithm for RM$(m,r)$ using subspaces of dimension $k$
@@
4:         Compute the projections $\mathbf{Y}_{/\mathbb{B}_{i}}$, $i\in[n_{k,m}]$.
5:         if $r\neq k+1$ then
6:              For each $i\in[n_{k,m}]$, set $\widehat{\mathbf{Y}}_{/\mathbb{B}_{i}}\leftarrow$ RPA($\mathbf{Y}_{/\mathbb{B}_{i}},m-k,r-k,M_{\text{iter}}$).
@@
8:              For each $i\in[n_{k,m}]$, decode $\mathbf{Y}_{/\mathbb{B}_{i}}$ to $\widehat{\mathbf{Y}}_{/\mathbb{B}_{i}}$ using the FHT-based decoder.
9:         Compute $\overline{\mathbf{Y}}\leftarrow$ Aggregate$(\mathbf{Y},\widehat{\mathbf{Y}}_{/\mathbb{B}_{1}},\ldots,\widehat{\mathbf{Y}}_{/\mathbb{B}_{n_{k,m}}})$.
@@
1:procedure Aggregate($\mathbf{Y},\widehat{\mathbf{Y}}_{/\mathbb{B}_{1}},\ldots,\widehat{\mathbf{Y}}_{/\mathbb{B}_{n_{k,m}}}$)
2:     Initialize the length-$2^{m}$ vector $\textsc{Flip}\leftarrow\mathbf{0}$.
@@
4:         Compute $\phi(\mathbf{x})=\sum_{i=1}^{n_{k,m}}\mathds{1}\{Y_{/\mathbb{B}_{i}}([\mathbf{x}+\mathbb{B}_{i}])\neq\widehat{Y}_{/\mathbb{B}_{i}}([\mathbf{x}+\mathbb{B}_{i}])\}$.
5:         Set Flip$(\mathbf{x})=1$, if $\phi(\mathbf{x})>\frac{n_{k,m}}{2}$.
@@ III Main Results
An important fact about the error probabilities above is [23, Prop. 2], which states that for all $m\geq 1,1\leq r\leq m$, we have that the probability of error over the BSC$(p)$, upon transmission of a codeword $\mathbf{c}\in\text{RM}(m,r)$ is independent of $\mathbf{c}$. This, in turn, implies that we can focus our analysis on the case when $\mathbf{c}=\mathbf{0}\in\{0,1\}^{N}$, with $P_{\text{err}}(\text{RM}(m,r))=P_{\text{err}}(\mathbf{0})$. In all the analysis that follows, we shall make the simplifying assumption that the input codeword is fixed to be the all-zeros codeword $\mathbf{0}$.
Before we state our main results, we will require some more notation. Let $\overline{p}:=\frac{1}{2}\cdot(1-(1-2p)^{2^{r-2}})$ and for any $\alpha\in(0,\frac{1}{2})$, let $\eta(\alpha):=\frac{1}{2}\cdot(1-4\alpha(1-\alpha))$. Note that $\eta(\alpha)<\frac{1}{2}$, for all $\alpha\in(0,\frac{1}{2})$. In what follows, we first focus on performance guarantees of RPA decoding when one-dimensional subspaces are used for projection. We then discuss some consequences of our main result on error probabailities using such subspaces and then state our result on error probabilities of RPA decoding using subspaces of dimension $k>1$. Our first main result is encapsulated in the following theorem:
@@
For any $0<\epsilon<\eta(\overline{p})$, we have that for $r\geq 2$, using one-dimensional subspaces for projection,
$P_{\text{err}}(\text{RM}(m,r))\leq 32N^{r+1}\cdot\text{exp}\left(-2^{-r-1}{N\epsilon^{2}}\right).$
@@
$\rho_{m}(r,\delta):=m(r+1)-\frac{\delta\log_{2}(e)}{8}\cdot 2^{m-r}\cdot(1-2p)^{2^{r}}+5.$
@@
Note, however, that RM codes with $r\leq\log(\overline{c}m)$, where $\overline{c}<c$, have rate $R_{m}:=\frac{\log|\text{RM}(m,r)|}{N}=\frac{{m\choose\leq r}}{N}$ such that $R_{m}\xrightarrow{m\to\infty}0$. In other words, via the results in this paper, we only obtain that RPA decoding achieves vanishing error probabilities, in the limit as the blocklength goes to infinity, for selected RM codes that have vanishing rate. An important question hence, as mentioned earlier, is whether RM codes can achieve positive rate (or even the capacity of the BSC$(p)$) under RPA decoding—a fertile topic for future research.
Now, for fixed (constant) orders $r$, Theorem III.1 also allows us to understand how quickly, as a function of $m$, the crossover probability $p$ (also called the “decoding radius") can converge to its maximum value that is $1/2$, while also ensuring that $P_{\text{err}}(\text{RM}(m,r))\xrightarrow{m\to\infty}0$. Observe that the statement that a decoding algorithm can achieve vanishing error probabilities, in the limit as the blocklength $N$ tends to infinity, is equivalent to the statement that the algorithm can correct $N(p-\delta)$ errors with high probability, for any $\delta\in(0,1)$. The following result hence establishes a lower bound on the number of such errors correctable with high probability, as a function of $m$, for fixed orders $r$, in a vein similar to the results in [15]. Recall that $d=2^{m-r}$ is the minimum distance of RM$(m,r)$.
@@
This implies that the convergence of the decoding radius to $1/2$ (captured by $\gamma_{m}$), for fixed orders $r\geq 2$, under our analysis of RPA decoding, is slower than that obtainable using the decoding algorithm $\Phi_{r}^{m}$ in [15] (captured by $\beta_{m}$). However, for sufficiently large $m$, it can be verified similarly that the convergence of the decoding radius to $1/2$ under RPA decoding is faster than the decoding algorithm $\Psi_{r}^{m}$ in [15]. It is an interesting question to check whether the RPA decoder can outperform $\Phi_{r}^{m}$ in [15] in terms of the speed of convergence of the decoding radius, using more sophisticated analysis of the RPA error probabilities, possibly by considering cases where each RPA run converges after two or more iterations.
Our final result obtains an upper bound on the probability of error of RPA decoding using subspaces of dimension $k>1$ for projection. For any $j\in[r-1]$, let $p^{(j)}:=\frac{1}{2}\cdot\left(1-(1-2p)^{2^{j}}\right)$; note that $\overline{p}=p^{(r-2)}$.
Theorem III.2.
For any $0<\epsilon<\eta(\overline{p})$, we have that for $r\geq 2$, using $k$-dimensional subspaces for projection, where $k|(r-1)$,
$P_{\text{err}}(\text{RM}(m,r))\leq 64N^{3}\cdot n_{k,m}^{\frac{r-1}{k}}\cdot\text{exp}\left(-\ln\left(\frac{1-p^{(r-k-1)}}{p^{(r-k-1)}}\right)\cdot 2^{-r-1-k}N\epsilon^{2}\right).$
We argue in Section VI, using Theorem III.2, that the guarantee on the growth rate of $r$ with $m$ to allow for vanishing error probabilities is the same as that in Corollary III.1 above (see Corollary VI.1 for details). Although one expects that the use of higher-dimensional subspaces for projection should lead to significantly improved estimates of error probabilities, our current analysis is unable to demonstrate such an improvement, mainly owing to the use of certain somewhat generic, albeit powerful, results to show the concentration of some random variables of interest about their expected value. We believe that strengthening this concentration inequality in the setting of higher-dimensional subspaces will lead to better bounds on the $P_{\text{err}}(\text{RM}(m,r))$—a task that we leave for future work.
III-A Organization of the Paper
Sections IV and V are dedicated to the setting where one-dimensional subspaces are used for projection, and lead to the proof of Theorem III.1. Theorem III.2 is then proved in Section VI, using the tools introduced in the previous sections.
In Section IV, we shall first analyze the behaviour of the FHT decoder in Step 8 of Algorithm 1 (which acts on the first-order RM codes obtained via the projection step) and the aggregation step in Step 9 of Algorithm 1, separately. Such analysis immediately leads to an understanding of the performance of the RPA decoder on second-order ($r=2$) RM codes. In Section V, we shall use the results obtained on the error probabilities of the FHT decoder and in the aggregation step, recursively, to get estimates of the error probabilities of the RPA decoder, using one-dimensional subspaces, for general orders $r>2$. With the recursive analysis of error probabilities in place in Section V, we then briefly discuss the modifications that need to be made to extend the analysis to the setting where higher-dimensional subspaces are used in the RPA algorithm, in Section VI, thus leading to a proof of Theorem III.2.